OpenStudy (anonymous):

If f(x) = 4x – 1 and g(x) = x2 + 2x + 2, what is f(g(x)) ? Answer: A)4x^2 + 8x + 7 B)x^2 + 8x – 2 C)4x^2 + 6x – 1 D)x^2 + 6x + 1 please help:)

OpenStudy (anonymous):

f(x) = 4x – 1 g(x) = x^2 + 2x + 2 f(g(x))=4(x^2 + 2x + 2) - 1 f(g(x))=4x^2 + 8x + 8 - 1 f(g(x))=4x^2 + 8x + 7 So, it's A. :)